OpenStudy (anonymous):

Students in Mr. Klein’s chemistry class are conducting an experiment to calculate the specific heat of a copper sample. Mr. Klein requires each lab team to repeat the experiment until their percent error is 2% or less. Percent error is calculated using observed values (those students calculate from their lab results) and the true value (the accepted value based on reference). The formula for percent error is:(observed – true) x 100 true If the accepted value for the specific heat of copper is 0.385 J/g°C, which of the following lab teams have an unresolved problem on their hands? a. team 1: calculates 0.392 J/gºC b. team 2: calculates 0.395 J/gºC c. both teams d. neither team

OpenStudy (anonymous):

As of right now i know for sure its not A or C.

OpenStudy (aaronq):

use the formula to test both possibilities: percent error = \(\dfrac{observed – true}{true} * 100\) Team1 1: percent error = \(\dfrac{| 0.392– 0.385|}{0.385} * 100\%=1.82\%\) So team 1 is good, their percent error is less than 2%. Now do team 2. (by the way, the formula you were given is wrong.)

OpenStudy (anonymous):

thank you aaronq

OpenStudy (aaronq):

no problem, dude!

OpenStudy (anonymous):

If I did this right team 2 should of came around 2.60% making them having a problem.

OpenStudy (aaronq):

thats correct \(\checkmark\)

OpenStudy (anonymous):

thank you once again

OpenStudy (aaronq):

no worries!